# BEND: Benchmarking DNA Language Models on Biologically Meaningful Tasks

### v1

@@ 2.1 Eukaryotic DNA organization and terminology
Figure 1: The organization of eukaryotic genomic DNA. The numbers are indicative examples for the human genome. Genes are structured as alternating introns (average: 5,400 bp) and exons (average: 170 bp), and have a promoter regulatory element before their transcription start site. Enhancer regulatory elements can be present multiple thousands bases away from the gene. DNA is wrapped around histone proteins (yellow cylinders) and densely packed as a chromosome.
@@ 2.3.1 DNA language models
GENA-LM (Fishman et al., 2023) proposed multiple medium-size LMs trained on human and multi-species genomes based on BERT and the BigBird (Zaheer et al., 2020) architecture for long sequences. Byte-Pair Encoding (BPE) was used for tokenization to further increase the receptive field, enabling an input length of about 36,000 bp. Models were evaluated on tasks comprising promoter, SS, enhancer, chromatin profile and polyadenylation site prediction. While covering the same biological phenomena, tasks were defined differently than in NT. Similarly, DNABERT-2 (Zhou et al., 2023) replaced DNABERT’s k-mer tokenizer with BPE and pre-trained on multi-species genomes.
@@ 2.3.3 Benchmark collections on DNA
Genomic Benchmarks (Grešová et al., 2023) features a collection of balanced classification tasks on DNA sequences with a median length ranging from 200 to 2,381 bp. The benchmark covers the classification of DNA elements and the prediction of a sequence’s origin. The element classification tasks are defined on human DNA, with one of the tasks covering D. melanogaster additionally. For each task, only performance of a baseline supervised neural network model was reported.
@@ 3 Tasks and Datasets
We introduce the collection of tasks included in BEND. For each task, we additionally provide a datasheet (Gebru et al., 2018) in section A.1. All tasks are provided in bed format, listing the genome coordinates of samples (A.2). This makes it convenient to include more flanking context without reprocessing the data, should future works find it useful to take more bp into account. Data splits were made by holding out whole chromosomes, with the exception of gene finding, where a sequence identity based split is recommended.
@@
Task Type (# labels) # Samples Length range Evaluation (# train/val/test) Metric Source Gene finding Nucleotide-wise Multiclass (9) 5,976 1,433 - 14,000 bp 4780/597/597 MCC GENCODE (Frankish et al., 2021) Enhancer annotation Binned (128bp) Binary 285 100,096 bp 10-fold CV AUPRC Fulco et al. (2019), Gasperini et al. (2019), Enformer (Avsec et al., 2021) Chromatin accessibility Sequence-wise Multilabel (125) 2,005,617 512 bp 1,354,042/ 279,422/372,153 AUROC ENCODE Project Consortium (2012) Histone modification Sequence-wise Multilabel (19) 612,081 512 bp 420,713/ 70,801/120,567 AUROC ENCODE Project Consortium (2012) CpG methylation Sequence-wise Multilabel (7) 959,039 512 bp 743,095/ 109,717/106,227 AUROC ENCODE Project Consortium (2012) Noncoding variant effects (expression) Sequence-wise Binary 106,221 512 bp zero-shot AUROC DeepSEA (Zhou & Troyanskaya, 2015) Noncoding variant effects (disease) Sequence-wise Binary 295,495 512 bp zero-shot AUROC ClinVar (Landrum et al., 2020)
@@ 3.2 Enhancer annotation
Metric  The AUPRC for binary classification is computed over all labels. As the number of samples is too limited for measuring performance robustly on a single test split, we perform 10-fold cross-validation in order to be able to evaluate performance over all samples.
@@ 3.3 Chromatin accessibility prediction
Data  DNase I hypersensitive sites were obtained from ENCODE (ENCODE Project Consortium, 2012; Luo et al., 2020; Kagda et al., 2023; Hitz et al., 2023) for 125 cell types. Following the preprocessing of Kelley et al. (Kelley et al., 2016), segments of length 512 bp were labeled with binary vectors $\mathbf{y}\in\{0,1\}^{125}$, with $y_{i}=1$ if the chromatin is open for the $i$’th cell type. Detailed processing is laid out in A.1.3. Samples were partitioned based on chromosomes.
@@ 4 Modeling
Table 2: Overview of the LMs included in the benchmark. Models from all works proposing DNA LMs applicable to the human genome were considered.
Model Seq length Trained on Architecture Source AWD-LSTM Infinitea Multispecies RNN This work Dilated ResNet 10,000 Human Refc CNN This work Nucleotide Transformer 5,994 Human Refc BERT Dalla-Torre et al. (2023) Nucleotide Transformer 5,994 Multispecies BERT Dalla-Torre et al. (2023) Nucleotide Transformer 5,994 1000 Genomes projectd BERT Dalla-Torre et al. (2023) Nucleotide Transformer V2 12,282 Multispecies BERT Dalla-Torre et al. (2023) DNABERT 512 Human Refc BERT Ji et al. (2021) DNABERT-2 Infiniteb Multispecies BERT Zhou et al. (2023) GENA-LM 4500 1000 Genomes projectd BERT Fishman et al. (2023) GENA-LM 36,000 1000 Genomes projectd BigBird Fishman et al. (2023) HyenaDNA 1,000,000 Human Refc Hyena Nguyen et al. (2023) HyenaDNA 1,000 Human Refc Hyena Nguyen et al. (2023) a As the LSTM compresses all preceding tokens into a single hidden state, it can technically process infinite sequences, even though it was trained at finite lengths and might not have learnt to exploit such long contexts. b DNABERT-2 uses ALiBi (Press et al., 2022) to encode position, which can technically scale to any sequence length. In practice, the model was trained on finite lengths and the authors recommend embedding sequences below 10,000 bp. c (Schneider et al., 2017), d (McVean et al., 2012)
@@
For each task, we train two supervised models without pre-training. For a direct comparison of raw and embedded DNA, we train the two-layer CNN on one-hot encoded sequences. For chromatin accessibility, histone modificaton and CpG methylation prediction, we train the Basset model (Kelley et al., 2016), which was specifically designed for modeling genome-wide functional genomics data. For gene finding and enhancer annotation, we train the ResNet CNN model on one-hot encoded sequences. For variant effect prediction, no supervised models are trained. For all tasks where Basset is not applicable, we also report the performance of a previously published task-specific non-LM expert method on the benchmark dataset to put LM performance into context.
@@ 5 Results
Gene finding Enhancer annotation Chromatin accessibility Histone modification CpG Methylation Variant effects (expression) Variant effects (disease) Expert method 0.80 Augustus 0.07 $\pm$ 0.04 Enformer 0.85 Basset 0.74 Basset 0.93 Basset 0.70 DeepSEA 0.56 DeepSEA Fully supervised ResNet 0.46 0.06 - - - - - CNN 0.00 0.03 0.75 0.76 0.84 - - Pre-trained ResNet-LM 0.36 0.02 0.82 0.77 0.87 0.56 0.55 AWD-LSTM 0.05 0.03 0.69 0.74 0.81 0.51 0.48 NT-H 0.41 0.05 0.74 0.76 0.88 0.55 0.48 NT-MS 0.68 0.06 0.79 0.78 0.92 0.55 0.77 NT-1000G 0.49 0.04 0.77 0.77 0.89 0.45 0.49 NT-V2 0.64 0.05 0.80 0.76 0.91 0.48 0.48 DNABERT 0.20 0.03 0.85 0.79 0.91 0.60 0.56 DNABERT-2 0.43 0.03 0.81 0.78 0.90 0.49 0.51 GENA-LM BERT 0.52 0.03 0.76 0.78 0.91 0.49 0.55 GENA-LM BigBird 0.39 0.04 0.82 0.78 0.91 0.49 0.52 HyenaDNA large 0.35 0.03 0.84 0.76 0.91 0.45 0.51 HyenaDNA tiny 0.10 0.02 0.78 0.76 0.86 0.47 0.44
@@
DNABERT shows the highest performance, on par with the specialized Basset model (0.85). All other LMs perform worse on this task, offering no advantage over the fully supervised expert model.
@@ 6 Discussion
We introduce BEND, a benchmark collection for DNA LMs that aims at better recapitulating the fundamental aspects of genomic sequence data. We find that currently available DNA LMs already show promising performance on some tasks over fully supervised baselines, but do not offer consistent improvements over all included tasks and can fall short of surpassing specialized existing prediction methods. Overall, we find that NT-MS is a strong default LM, but is in some tasks inferior to the much smaller DNABERT. Interestingly, while both models trained using the MLM objective, we find that they learned distinct genomic features during pre-training. With the pre-training data and the tokenization strategy being the key architectural difference, these choices may deserve more attention in future DNA LMs.
@@
On the gene finding task, we observe that the combination of NT-MS with a simple two-layer CNN shows a very promising performance level compared to the specialised AUGUSTUS, which was found to be the state of the art in a recent benchmark (Scalzitti et al., 2020). This suggests that future more sophisticated LM-based gene finders might become a method of choice for this problem. The result also indicates that current DNA LMs are capable of modeling long-range dependencies to some extent.
@@ 7 Limitations and Outlook
Acknowledgements and disclosure of funding
This work was funded in part by Innovation Fund Denmark (0153-00197B), the Novo Nordisk Foundation through the MLLS Center (Basic Machine Learning Research in Life Science, NNF20OC0062606), the Pioneer Centre for AI (DRNF grant number P1), and the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516).
This work was supported by the Helmholtz Association under the joint research school ”Munich School for Data Science (MUDS)”.
We would like to acknowledge and thank Ziga Avsec, David Kelley, as well as the rest of the authors behind the Enformer model Avsec et al. (2021), for providing the set of transcription start sites used in the enhancer annotation task.
@@ A.1.6 Noncoding variant effects (Expression)
  • Preprocessing As the original genomic coordinates refer to the previous reference genome GRCh37, we used LiftOver to transfer the coordinates to the current reference GRCh38. Any coordinates that could not be mapped were discarded. We only use SNPs included in fold 0. We applied Ensembl VEP (McLaren et al., 2016) to categorize variants by consequence. VEP infers the consequence of a variant by comparing a variant’s position to the reference genome annotation, determining what type of sequence region (Table A4) the variant lies in. To obtain one consequence per variant, we use VEP’s --most_severe flag, returning the consequence with the potentially most severe effect on function. In DeepSEA, the adjacent 1,000 bp served as context for classification. As this exceeds the maximum context length of some of the benchmarked models, and chunking inputs is not a meaningful strategy when adjacent bps serve only as context for an unsupervised embedding, we use 512 bp instead. As this is a zero-shot task, no split is performed, with the full dataset serving as test set.
@@ A.1.7 Noncoding variant effects (Disease)
Intron variant 56,183 5,046 8.24%
Intergenic variant 22,622 757 3.24%
Upstream gene variant 5,849 584 9.08%
Downstream gene variant 4,206 439 9.45%
Regulatory region variant 3,800 249 6.15%
Noncoding transcript exon variant 2,797 344 10.95%
3’ UTR variant 1,621 428 20.89%
5’ UTR variant 415 55 11.70%
TF binding site variant 413 23 5.28%
Splice region variant 101 30 22.90%
splice polypyrimidine tract variant 87 18 17.14%
Missense variant 53 9 14.52%
@@ A.2 Formatting
Code to extract DNA sequences from the reference genome with the bed coordinates, dataloaders, models and config files is available on Github (https://github.com/frederikkemarin/BEND).
@@ A.4 Distribution
All data is available at https://sid.erda.dk/sharelink/f6hdp1zTzh or https://zenodo.org/records/10039745. Code, configs and scripts to extract data and run all experiments are provided at https://github.com/frederikkemarin/BEND.
@@ A.6.3 Upsampling of embeddings
LMs that make use of k-mer or byte-pair encoding (BPE) tokenization strategies return less embedding vectors than their original input sequence length. For nucleotide-level prediction tasks, an embedding sequence of equal length to the nucleotide-wise label sequence is needed. In order to benchmark all LMs equally, regardless of how they tokenize inputs, we upsample embeddings. For the 6-mer tokenization employed by NT, we repeat each embedding vector 6 times. For BPE in GENA-LM and DNABERT-2, we repeat each token’s embedding by the length of the token’s sequence. DNABERT, which uses overlapping k-mers, returns a reduced number of embeddings due to the fact that at the left and right borders of the sequence there is no k/2 context available to construct a k-mer embedding around the nucleotide. As DNABERT does not perform any padding to correct for this, these initial and terminal k-mer embeddings are missing. We repeat the first and the last embedding to match the original input sequence length. For k=6, we repeat the first embedding two and the last embedding three times.
@@ A.8 Extended results
CDSF (0) DonorF (1) IntronF (2) AcceptorF (3) CDSR (4) AcceptorR (5) IntronR (6) DonorR (7) Intergenic (8) Model Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision AUGUSTUS 0.89 0.90 0.80 0.88 0.83 0.87 0.79 0.86 0.89 0.91 0.81 0.86 0.85 0.89 0.80 0.85 0.86 0.81 ResNet 0.79 0.84 0.7 0.81 0.43 0.59 0.63 0.78 0.84 0.83 0.74 0.83 0.68 0.52 0.61 0.76 0.59 0.62 CNN 0.0 0.0 0.0 0.0 0.01 0.28 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.0 0.39 AWD-LSTM 0.0 0.33 0.0 0.0 0.26 0.28 0.0 0.0 0.0 0.22 0.0 0.0 0.02 0.38 0.0 0.0 0.81 0.4 ResNet-LM 0.59 0.62 0.0 0.0 0.52 0.41 0.0 0.0 0.51 0.76 0.0 0.0 0.56 0.46 0.0 0.0 0.43 0.55 NT-H 0.67 0.59 0.0 0.0 0.59 0.49 0.0 0.0 0.6 0.7 0.0 0.22 0.65 0.54 0.0 0.0 0.49 0.66 NT-MS 0.94 0.89 0.73 0.66 0.84 0.69 0.5 0.73 0.93 0.89 0.64 0.74 0.86 0.69 0.57 0.66 0.55 0.79 NT-1000G 0.78 0.79 0.03 0.28 0.7 0.59 0.01 0.64 0.76 0.84 0.14 0.62 0.74 0.63 0.06 0.43 0.57 0.7 NT-V2 0.94 0.91 0.75 0.73 0.78 0.65 0.55 0.8 0.94 0.91 0.75 0.74 0.81 0.68 0.59 0.77 0.57 0.77 DNABERT 0.43 0.49 0.47 0.33 0.54 0.34 0.24 0.36 0.39 0.58 0.2 0.52 0.2 0.45 0.38 0.35 0.56 0.5 DNABERT2 0.51 0.69 0.09 0.42 0.6 0.5 0.0 0.0 0.51 0.69 0.13 0.49 0.57 0.56 0.0 0.0 0.63 0.65 GENA-LM BERT 0.82 0.81 0.34 0.83 0.69 0.6 0.29 0.59 0.82 0.81 0.26 0.57 0.7 0.61 0.31 0.65 0.53 0.65 GENA-LM BigBird 0.41 0.53 0.13 0.35 0.59 0.49 0.13 0.33 0.39 0.56 0.11 0.33 0.75 0.51 0.04 0.41 0.43 0.66 HyenaDNA tiny 0.17 0.26 0.05 0.11 0.29 0.33 0.02 0.23 0.02 0.39 0.06 0.43 0.04 0.46 0.0 0.0 0.79 0.41 HyenaDNA large 0.23 0.4 0.04 0.18 0.6 0.45 0.06 0.23 0.36 0.4 0.23 0.38 0.62 0.52 0.0 0.08 0.48 0.62

### v3

@@ 2.1 Eukaryotic DNA organization and terminology
Figure 1: The organization of eukaryotic genomic DNA. Numbers are indicative examples for the human genome. Genes are structured as alternating introns (average: 5,400 bp) and exons (average: 170 bp), and have a promoter regulatory element before their TSS. Enhancers can be thousands of bp away from the gene. DNA is wrapped around histone proteins and densely packed as a chromosome.
@@ 2.3.1 DNA language models
GENA-LM (Fishman et al., 2023) proposed multiple medium-size LMs trained on human and multi-species genomes based on BERT and the BigBird (Zaheer et al., 2020) architecture for long sequences. Byte-Pair Encoding (BPE) was used for tokenization to further increase the receptive field, enabling an input length of about 36,000 bp. Models were evaluated on tasks comprising promoter, SS, enhancer, chromatin profile and polyadenylation site prediction. While covering the same biological phenomena, tasks were defined differently than in NT. Similarly, DNABERT-2 (Zhou et al., 2023) replaced DNABERT’s k-mer tokenizer with BPE and pre-trained on multi-species genomes, while GROVER (Sanabria et al., 2023) adopted BPE for the human genome.
@@ 2.3.3 Benchmark collections on DNA
Genomic Benchmarks (Grešová et al., 2023) features balanced classification tasks on DNA sequences with median lengths ranging from 200 to 2,381 bp. The benchmark covers the classification of functional elements and the prediction of a sequence’s origin. The element classification tasks are defined on human DNA, with one task covering D. melanogaster additionally. For each task, only performance of a baseline supervised neural network model was reported.
@@ 3 Tasks and Datasets
We introduce the collection of tasks included in BEND. For each task, we additionally provide a datasheet (Gebru et al., 2018) in section A.1. All tasks are provided in bed format, listing the genome coordinates of samples (A.2). This makes it convenient to include more flanking context without reprocessing the data, should future works find it useful to take more bp into account.
@@
Task Type (# labels) # Samples Length range Evaluation (# train/val/test) Metric Source Gene finding Nucleotide-wise Multiclass (9) 5,976 1,433 - 14,000 bp 4780/597/597 MCC GENCODE (Frankish et al., 2021) Enhancer annotation Binned (128bp) Binary 285 100,096 bp 10-fold CV AUPRC Fulco et al. (2019), Gasperini et al. (2019), Enformer (Avsec et al., 2021) Chromatin accessibility Sequence-wise Multilabel (125) 2,005,617 512 bp 1,354,042/ 279,422/372,153 AUROC ENCODE Project Consortium (2012) Histone modification Sequence-wise Multilabel (19) 612,081 512 bp 420,713/ 70,801/120,567 AUROC ENCODE Project Consortium (2012) CpG methylation Sequence-wise Multilabel (7) 959,039 512 bp 743,095/ 109,717/106,227 AUROC ENCODE Project Consortium (2012) Noncoding variant effects (expression) Sequence-wise Binary 105,263 512 bp zero-shot AUROC DeepSEA (Zhou & Troyanskaya, 2015) Noncoding variant effects (disease) Sequence-wise Binary 295,495 512 bp zero-shot AUROC ClinVar (Landrum et al., 2020)
@@ 3.2 Enhancer annotation
Metric  The AUPRC is computed over all labels. As the number of samples is too limited for measuring performance robustly on a single test split, we perform 10-fold cross-validation in order to evaluate performance over all samples.
@@ 3.3 Chromatin accessibility prediction
Data  DNase I hypersensitive sites were obtained from ENCODE (ENCODE Project Consortium, 2012) for 125 cell types. Following the preprocessing of Kelley et al. (2016), segments of length 512 bp were labeled with binary vectors $\mathbf{y}\in\{0,1\}^{125}$, with $y_{i}=1$ if the chromatin is open for the $i$’th cell type. Detailed processing is laid out in A.1.3. Samples were partitioned based on chromosomes.
@@ 4 Modeling
Table 2: Overview of the LMs applicable to the human genome included in the benchmark.
Model Seq length Trained on Architecture Source AWD-LSTM Infinitea Multispecies RNN This work Dilated ResNet 10,000 Human Refd CNN This work Nucleotide Transformer 5,994 Human Refd BERT Dalla-Torre et al. (2023) Nucleotide Transformer 5,994 Multispecies BERT Dalla-Torre et al. (2023) Nucleotide Transformer 5,994 1000 Genomes projectd BERT Dalla-Torre et al. (2023) Nucleotide Transformer V2 12,282 Multispecies BERT Dalla-Torre et al. (2023) DNABERT 512 Human Refd BERT Ji et al. (2021) DNABERT-2 Infiniteb Multispecies BERT Zhou et al. (2023) GENA-LM 4,500 1000 Genomes projecte BERT Fishman et al. (2023) GENA-LM 36,000 1000 Genomes projecte BigBird Fishman et al. (2023) HyenaDNA 1,000,000 Human Refd Hyena Nguyen et al. (2023) HyenaDNA 1,000 Human Refd Hyena Nguyen et al. (2023) GROVER 8,160c Human Refd BERT Sanabria et al. (2023) a As the LSTM compresses all preceding tokens into a single hidden state, it can technically process infinite sequences, even though it was trained at finite lengths and might not have learnt to exploit such long contexts. b DNABERT-2 uses ALiBi (Press et al., 2022) to encode position, which can technically scale to any sequence length. In practice, the model was trained on finite lengths and the authors recommend embedding sequences below 10,000 bp. c No explicit length was reported in bp. The indicated number was derived by considering 510 BPE tokens of size 16. d Schneider et al. (2017), e McVean et al. (2012)
@@
For each task, we train two supervised models without pre-training. For a direct comparison of raw and embedded DNA, we train the two-layer CNN on one-hot encoded sequences. For chromatin accessibility, histone modificaton and CpG methylation prediction, we train the Basset model (Kelley et al., 2016), which was specifically designed for modeling genome-wide functional genomics data. For gene finding and enhancer annotation, we train the ResNet CNN model on one-hot encoded sequences. For variant effect prediction, no supervised models are trained. For all tasks where Basset is not applicable, we report the performance of a previously published task-specific expert method on the benchmark dataset to put LM performance into context.
@@ 5 Results
Gene finding Enhancer annotation Chromatin accessibility Histone modification CpG Methylation Variant effects (expression) Variant effects (disease) Expert method 0.80 Augustus 0.07 Enformer 0.85 Basset 0.74 Basset 0.93 Basset 0.70 DeepSEA 0.56 DeepSEA Fully supervised ResNet 0.46 0.06 - - - - - CNN 0.00 0.03 0.75 0.76 0.84 - - Pre-trained ResNet-LM 0.36 0.02 0.82 0.77 0.87 0.55 0.55 AWD-LSTM 0.05 0.03 0.69 0.74 0.81 0.53 0.45 NT-H 0.41 0.05 0.74 0.76 0.88 0.55 0.48 NT-MS 0.68 0.06 0.79 0.78 0.92 0.54 0.77 NT-1000G 0.49 0.04 0.77 0.77 0.89 0.45 0.49 NT-V2 0.64 0.05 0.80 0.76 0.91 0.48 0.48 DNABERT 0.20 0.03 0.85 0.79 0.91 0.60 0.56 DNABERT-2 0.43 0.03 0.81 0.78 0.90 0.49 0.51 GENA-LM BERT 0.52 0.03 0.76 0.78 0.91 0.49 0.55 GENA-LM BigBird 0.39 0.04 0.82 0.78 0.91 0.49 0.52 HyenaDNA large 0.35 0.03 0.84 0.76 0.91 0.51 0.45 HyenaDNA tiny 0.10 0.02 0.78 0.76 0.86 0.47 0.44 GROVER 0.28 0.03 0.82 0.77 0.89 0.56 0.51
@@
DNABERT shows the highest performance, on par with the specialized Basset model (0.85). All other LMs perform worse on this task, offering no advantage over Basset.
@@ 6 Discussion
We find that currently available DNA LMs already show promising performance on some tasks over fully supervised baselines, but do not offer consistent improvements over all included tasks and can fall short of surpassing specialized existing prediction methods. Overall, we find that NT-MS is a strong default LM, but is in some tasks inferior to the much smaller DNABERT. Interestingly, while both models trained using the MLM objective, we find that they learned distinct genomic features during pre-training. With the pre-training data and the tokenization strategy being the key architectural difference, these choices may deserve more attention in future DNA LMs.
@@
On the gene finding task, we observe that NT-MS with a simple two-layer CNN shows promising performance compared to the specialized AUGUSTUS, which was found to be the state of the art in a recent benchmark (Scalzitti et al., 2020). This suggests that future more sophisticated LM-based gene finders might become a method of choice for this problem. The result also indicates that current DNA LMs are capable of modeling long-range dependencies to some extent.
@@ A.1.6 Noncoding variant effects (Expression)
  • Preprocessing As the original genomic coordinates refer to the previous reference genome GRCh37, we used LiftOver to transfer the coordinates to the current reference GRCh38. Any coordinates that could not be mapped were discarded. Variants where the original reference nucleotide does not match the nucleotide at the indicated position in GRCh38 were removed. We only use SNPs included in fold 0. We applied Ensembl VEP (McLaren et al., 2016) to categorize variants by consequence. VEP infers the consequence of a variant by comparing a variant’s position to the reference genome annotation, determining what type of sequence region (Table A4) the variant lies in. To obtain one consequence per variant, we use VEP’s --most_severe flag, returning the consequence with the potentially most severe effect on function. In DeepSEA, the adjacent 1,000 bp served as context for classification. As this exceeds the maximum context length of some of the benchmarked models, and chunking inputs is not a meaningful strategy when adjacent bps serve only as context for an unsupervised embedding, we use 512 bp instead. As this is a zero-shot task, no split is performed, with the full dataset serving as test set.
@@ A.1.7 Noncoding variant effects (Disease)
Intron variant 55,710 5,002 8.24%
Intergenic variant 22,465 753 3.24%
Upstream gene variant 5,760 579 9.13%
Downstream gene variant 4,146 435 9.50%
Regulatory region variant 3,762 248 6.18%
Noncoding transcript exon variant 2,757 342 11.03%
3’ UTR variant 1,599 426 21.03%
5’ UTR variant 408 54 11.69%
TF binding site variant 410 23 5.31%
Splice region variant 99 30 23.26%
splice polypyrimidine tract variant 85 18 17.48%
Missense variant 51 9 15.00%
@@ A.2 Formatting
Code to extract DNA sequences from the reference genome with the bed coordinates, dataloaders, models and config files is available on Github (https://anonymous.4open.science/r/BEND-8C42/README.md).
@@ A.4 Distribution
All data is available at https://sid.erda.dk/cgi-sid/ls.py?share_id=aNQa0Oz2lY Code, configs and scripts to extract data and run all experiments are provided at https://github.com/frederikkemarin/BEND.
@@ A.6.3 Upsampling of embeddings
LMs that make use of k-mer or byte-pair encoding (BPE) tokenization strategies return less embedding vectors than their original input sequence length. For nucleotide-level prediction tasks, an embedding sequence of equal length to the nucleotide-wise label sequence is needed. In order to benchmark all LMs equally, regardless of how they tokenize inputs, we upsample embeddings. For the 6-mer tokenization employed by NT, we repeat each embedding vector 6 times. For BPE in GENA-LM, DNABERT-2 and GROVER, we repeat each token’s embedding by the length of the token’s sequence. DNABERT, which uses overlapping k-mers, returns a reduced number of embeddings due to the fact that at the left and right borders of the sequence there is no k/2 context available to construct a k-mer embedding around the nucleotide. As DNABERT does not perform any padding to correct for this, these initial and terminal k-mer embeddings are missing. We repeat the first and the last embedding to match the original input sequence length. For k=6, we repeat the first embedding two and the last embedding three times.
@@ A.8 Extended results
CDSF (0) DonorF (1) IntronF (2) AcceptorF (3) CDSR (4) AcceptorR (5) IntronR (6) DonorR (7) Intergenic (8) Model Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision Recall Precision AUGUSTUS 0.89 0.90 0.80 0.88 0.83 0.87 0.79 0.86 0.89 0.91 0.81 0.86 0.85 0.89 0.80 0.85 0.86 0.81 ResNet 0.79 0.84 0.7 0.81 0.43 0.59 0.63 0.78 0.84 0.83 0.74 0.83 0.68 0.52 0.61 0.76 0.59 0.62 CNN 0.0 0.0 0.0 0.0 0.01 0.28 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 1.0 0.39 AWD-LSTM 0.0 0.33 0.0 0.0 0.26 0.28 0.0 0.0 0.0 0.22 0.0 0.0 0.02 0.38 0.0 0.0 0.81 0.4 ResNet-LM 0.59 0.62 0.0 0.0 0.52 0.41 0.0 0.0 0.51 0.76 0.0 0.0 0.56 0.46 0.0 0.0 0.43 0.55 NT-H 0.67 0.59 0.0 0.0 0.59 0.49 0.0 0.0 0.6 0.7 0.0 0.22 0.65 0.54 0.0 0.0 0.49 0.66 NT-MS 0.94 0.89 0.73 0.66 0.84 0.69 0.5 0.73 0.93 0.89 0.64 0.74 0.86 0.69 0.57 0.66 0.55 0.79 NT-1000G 0.78 0.79 0.03 0.28 0.7 0.59 0.01 0.64 0.76 0.84 0.14 0.62 0.74 0.63 0.06 0.43 0.57 0.7 NT-V2 0.94 0.91 0.75 0.73 0.78 0.65 0.55 0.8 0.94 0.91 0.75 0.74 0.81 0.68 0.59 0.77 0.57 0.77 DNABERT 0.43 0.49 0.47 0.33 0.54 0.34 0.24 0.36 0.39 0.58 0.2 0.52 0.2 0.45 0.38 0.35 0.56 0.5 DNABERT2 0.51 0.69 0.09 0.42 0.6 0.5 0.0 0.0 0.51 0.69 0.13 0.49 0.57 0.56 0.0 0.0 0.63 0.65 GENA-LM BERT 0.82 0.81 0.34 0.83 0.69 0.6 0.29 0.59 0.82 0.81 0.26 0.57 0.7 0.61 0.31 0.65 0.53 0.65 GENA-LM BigBird 0.41 0.53 0.13 0.35 0.59 0.49 0.13 0.33 0.39 0.56 0.11 0.33 0.75 0.51 0.04 0.41 0.43 0.66 HyenaDNA tiny 0.17 0.26 0.05 0.11 0.29 0.33 0.02 0.23 0.02 0.39 0.06 0.43 0.04 0.46 0.0 0.0 0.79 0.41 HyenaDNA large 0.23 0.4 0.04 0.18 0.6 0.45 0.06 0.23 0.36 0.4 0.23 0.38 0.62 0.52 0.0 0.08 0.48 0.62 GROVER 0.31 0.53 0.12 0.29 0.55 0.39 0.01 0.14 0.45 0.48 0.26 0.36 0.47 0.48 0.06 0.25 0.48 0.55